# LoPA: Scaling dLLM Inference via Lookahead Parallel Decoding

### v1

@@ 4.2 Main Results
D2F-Dream-Base LoPA-Dist-NV S1 358.49 661.48 616.35 53.40 287.83 589.70 560.42 75.66
S2 413.79 1102.42 723.61 53.20 344.37 785.16 701.64 74.00
S3 475.83 1420.31 860.80 52.80 379.95 844.01 765.82 70.81
S4 503.30 1426.83 898.32 47.00 413.79 891.58 804.73 68.99
S5 547.92 1408.40 963.67 47.00 441.42 932.31 888.53 67.40
S6 568.50 1496.79 993.30 42.00 467.31 1074.57 971.05 63.61
S7 410.71 724.76 670.19 51.20 364.23 1042.55 771.30 70.58
S8 477.94 1286.34 832.42 50.00 423.83 973.03 830.27 70.51
S9 527.56 1387.88 854.29 46.60 463.85 1021.76 959.64 68.76
S10 573.79 1278.68 898.92 46.40 502.76 1114.08 977.91 65.88
S11 588.91 1349.33 921.64 44.80 525.66 1079.42 980.66 62.85
S12 630.28 1472.37 1063.34 40.60 566.97 1305.86 1079.47 60.12
@@
D2F-Dream-Instruct LoPA-Dist-NV S1 263.02 978.96 599.64 51.20 288.00 675.86 622.61 78.24
S2 315.50 1078.19 785.63 51.00 327.67 807.99 679.39 74.53
S3 378.42 1108.45 895.57 50.80 383.38 860.20 765.74 74.45
S4 426.59 1130.15 1017.69 47.00 419.70 979.73 849.78 72.48
S5 470.72 1550.00 1199.15 42.80 457.29 1015.40 921.06 68.39
S6 513.30 1656.67 1353.23 38.60 497.19 1403.30 1127.95 64.29
S7 281.80 715.75 553.48 51.80 329.86 770.33 691.63 77.63
S8 359.69 1105.22 783.33 51.40 393.16 778.01 742.47 74.68
S9 410.49 1311.42 941.31 50.80 437.08 1014.33 819.63 72.93
S10 462.90 1282.73 1040.49 47.20 469.80 1047.98 881.06 71.87
S11 506.30 1552.87 1159.76 41.80 506.10 1050.33 899.67 69.37
S12 543.32 1531.64 1332.85 37.00 536.71 1141.71 1017.81 65.50
@@
System Throughput. To fully unleash the parallelism enabled by LoPA across diverse hardware backends, we evaluate the wall-clock performance using our co-designed Branch Parallel (BP) inference system, as illustrated in Figure 3. Under multi-device deployment, the system exhibits near-linear scalability with respect to the branch count, effectively translating high TPF into tangible throughput gains. As visualized in Figure LABEL:fig:tps_comparison, the system achieves a single-sample throughput of 1073.86 tokens per second, where both the D2F baseline and D2F + LoPA employ identical, slightly lowered decoding thresholds to prioritize and maximize inference speed.

### v2

@@ 4.2 Main Results
D2F-Dream-Base LoPA-Dist-NV S1 415.19 813.04 720.35 53.00 345.52 959.05 704.39 75.97
S2 500.33 1185.77 874.87 53.40 402.52 913.12 842.83 73.54
S3 550.37 1472.41 929.72 51.20 436.22 994.82 885.27 71.19
S4 589.22 1576.93 1006.57 47.20 475.58 1203.61 1028.15 68.16
S5 633.16 1408.40 963.67 46.80 516.85 1212.65 1055.08 66.79
S6 678.26 1615.30 1150.65 41.80 546.72 1225.21 1121.57 64.14
S7 466.27 784.33 764.52 51.80 416.91 909.82 841.95 71.27
S8 545.90 1497.22 927.67 51.40 486.94 1176.14 959.37 68.39
S9 588.00 1584.28 983.09 48.60 520.70 1250.67 1056.01 68.01
S10 637.38 1552.56 1028.97 47.00 558.01 1115.26 1071.66 65.05
S11 655.45 1535.10 1059.72 43.80 592.94 1315.93 1155.11 64.44
S12 708.48 1470.95 1132.78 39.80 619.33 1299.25 1201.18 60.88
@@
D2F-Dream-Instruct LoPA-Dist-NV S1 305.74 959.00 695.88 52.80 330.62 758.34 674.53 78.17
S2 373.23 1302.99 877.12 51.40 402.63 961.29 804.31 74.22
S3 451.62 1419.09 1143.30 53.00 444.73 943.22 870.85 73.39
S4 503.71 1779.60 1226.72 46.60 495.93 1131.64 941.23 72.48
S5 568.65 1660.89 1317.38 42.00 540.76 1185.14 1033.60 68.99
S6 615.95 1951.86 1542.82 37.60 568.75 1352.22 1139.06 65.88
S7 325.15 697.49 620.42 50.80 379.42 839.65 710.10 75.28
S8 408.37 1182.69 866.90 51.00 449.56 934.55 838.35 75.13
S9 465.55 1097.40 1016.91 50.60 497.47 1172.31 946.98 74.75
S10 544.72 1542.99 1145.55 46.80 539.28 1147.95 1021.96 71.34
S11 591.57 1578.00 1204.05 42.20 580.04 1292.18 1132.19 66.94
S12 636.55 1811.71 1500.59 36.00 609.90 1407.56 1159.28 65.50
@@
System Throughput. To fully unleash the parallelism enabled by LoPA across diverse hardware backends, we evaluate the wall-clock performance using our co-designed Branch Parallel (BP) inference system, as illustrated in Figure 3. Under multi-device deployment, the system exhibits near-linear scalability with respect to the branch count, effectively translating high TPF into tangible throughput gains. As shown in Figure LABEL:fig:tps_comparison and Table 3, the system achieves a single-sample throughput of 1073.86 tokens per second, where both the D2F baseline and D2F + LoPA employ identical, slightly lowered decoding thresholds to prioritize and maximize inference speed.